OpenStudy (anonymous):

which equation shows how to calculate how many grams (g) of kcl would be produced from 4 mol koh the balanced reaction is mgcl2 + 2koh -> mg(oh)2 + 2kcl

OpenStudy (anonymous):

These are the answer choices

OpenStudy (anonymous):

Someone please help :(

OpenStudy (frostbite):

Hey @_97alex and welcome to Openstudy. The answers are a bit unmanageable but I'll do my best to explain: We know we got 4 mol of KOH which will react with MgCl2 and we assume that the reaction will run so that there is no KOH left. This accounts for the expression: \[\large \sf \frac{ 4~mol~KOH }{ 1 }\] Note that 2 mol KOH react to produce 2 KCl mol. This is why they have the expression: \[\large \sf \frac{ 2 ~mol~KCl }{ 2~mol~KOH }\] Lastly they multiply with molar mass, which they write as: \[\large \sf \frac{ 74.55~g~KCl }{ 1~mol }\] By doing these fractions, they make the expression way more difficult to look at, as we for example we usually just write the molar mass \(\sf 74.55~g/mol\) and not divide by 1 when we write the amount of substance of KOH.

OpenStudy (anonymous):

thank you!

OpenStudy (frostbite):

No problem at all. If you have further questions please let me know. :)